OpenStudy (anonymous):

Am I doing this correct? Let f(x) = 24/3x-2 Find f(-2).

OpenStudy (anonymous):

f(-2) = 24/3(-2) - 2 f(-2) = 24/-6 - 2 f(-2) = 24/-8 f(-2) = -3 so is the answer -3?

OpenStudy (anonymous):

brb I have to do something just let me know if it's correct or not and when i can come back you can explain if it's incorrect

OpenStudy (anonymous):

@sangya21 so when you gave me a medal does that mean i got it correct

OpenStudy (anonymous):

yups

OpenStudy (anonymous):

thank you so much!

OpenStudy (anonymous):

anytime :)